OpenStudy (anonymous):

Which shows 522 - 482 being evaluated using the difference of perfect squares method? (1 point) 522 - 482 = (2704 + 2304)(2704 - 2304) = 2,003,200 522 - 482 = (52 + 48)(52 - 48) = (100)(4) = 400 522 - 482 = 2704 - 2304 = 400 522 - 482 = (52 - 48)2 = (4)2 = 16

OpenStudy (anonymous):

@mathmath333 @MathNinja_C @sleepyjess

OpenStudy (anonymous):

none of the above

OpenStudy (anonymous):

... not possible

OpenStudy (anonymous):

522 is not a perfect square

OpenStudy (anonymous):

oops. suppose to be 52^2

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

and 48^2

OpenStudy (anonymous):

caps T_T

OpenStudy (anonymous):

\[(52^2-48^2)9=(52+48)(52-48)=400\]

OpenStudy (anonymous):

Thankss

OpenStudy (anonymous):

which is actually B, not C

OpenStudy (anonymous):

I stand corrected